### v2

# Hear the Scene: Audio-Enhanced Text Spotting

Abstract
Recent advancements in scene text spotting have focused on end-to-end methodologies that heavily rely on precise location annotations, which are often costly and labor-intensive to procure. In this study, we introduce an innovative approach that leverages only transcription annotations for training text spotting models, substantially reducing the dependency on elaborate annotation processes. Our methodology employs a query-based paradigm that facilitates the learning of implicit location features through the interaction between text queries and image embeddings. These features are later refined during the text recognition phase using an attention activation map. Addressing the challenges associated with training a weakly-supervised model from scratch, we implement a circular curriculum learning strategy to enhance model convergence. Additionally, we introduce a coarse-to-fine cross-attention localization mechanism for more accurate text instance localization. Notably, our framework supports audio-based annotation, which significantly diminishes annotation time and provides an inclusive alternative for individuals with disabilities. Our approach achieves competitive performance against existing benchmarks, demonstrating that high accuracy in text spotting can be attained without extensive location annotations.
Figure 1: Different annotation styles for text spotting. The blue point and lines are the position annotation for the text. The transcription annotation “Hills” is displayed in the top left corner of each image.
1 Introduction
Scene text spotting Tang et al. (2022a); Liu et al. (2020a); Xing et al. (2019); Liu et al. (2023, 2020b); Tang et al. (2022c, b) remains a vibrant area of research due to its extensive applications across various domains such as image-based translation, multimedia content analysis, and accessibility features in visual media. The field Liao et al. (2020); Zhao et al. (2024b); feng2023docpedia; Feng et al. (2023) has evolved from focusing primarily on horizontal or multi-oriented text to addressing more complex arbitrarily-shaped text. This evolution has prompted shifts from traditional annotation formats like horizontal rectangles and quadrilaterals to more sophisticated polygonal annotations, as depicted in Figure 1. Traditionally, scene text spotting involved separate processes for text detection and recognition, where both tasks required annotations for training—text recognition used transcription annotations, and text detection depended on geometric annotations like polygons or bounding boxes. The reliance on detailed geometric annotations for training not only escalates the cost but also increases the manpower required, making the process less efficient and scalable. Recent advancements have introduced methods Li et al. (2024b, a); Zhao et al. (2024c, a); Tang et al. (2024b, 2023, a) that directly identify and classify characters or words within texts, which, while reducing some dependencies, still rely heavily on character-level annotations and thus remain costly. Furthermore, the granularity of these annotations makes the task time-consuming and limits the feasibility of large-scale applications on real-world data. For instance, the annotation of text in videos or cluttered urban scenes can be prohibitively expensive and labor-intensive due to the complexity and volume of data. Given these challenges, there is a growing interest in exploring methods that reduce reliance on expensive annotations. Preliminary efforts, such as the single-point annotation strategy proposed by SPTS Liu et al. (2023), have marked a significant step forward but still leave room for improvement. This paper introduces a novel approach that rethinks scene text spotting from an image captioning perspective, where the text’s location is inferred implicitly through visual attention mechanisms, without direct location supervision. We propose a fully end-to-end architecture, termed EchoSpot, which leverages a query-based interaction between text queries and image embeddings to approximate text locations while ensuring robust text recognition. Our contributions are geared towards simplifying the annotation process, reducing costs, and making text spotting more accessible and scalable. By demonstrating that accurate text spotting can be achieved without explicit location annotations and by integrating novel strategies like circular curriculum learning and audio annotations, this work paves the way for future research and practical applications in the field of optical character recognition and beyond.
Figure 2: Overview of the proposed EchoSpot model architecture. The visual and contextual embeddings are first extracted by a backbone network. The embeddings are then decoded to focus on the relevant text regions using a query-based cross-attention module. Next, the refine-stage text query, the mask generated by the attention activation map, and image embeddings are decoded together to obtain a refined text position.
2 Related Work
The field of scene text spotting has witnessed a gradual shift from reliance on heavily annotated data towards methods that require fewer annotations. This evolution reflects the increasing need for cost-effective and scalable solutions. This section reviews existing literature, categorizing the studies into fully-supervised and weakly-supervised methods.
2.1 Fully-Supervised Text Spotting
Traditionally, text spotting techniques have drawn heavily from generic object detection frameworks, employing two-stage models that first detect text and then recognize it. Early approaches, like the one proposed by Li et al. Shi et al. (2016), utilized CNN-based detectors coupled with RNN-based recognition branches, applying RoI pooling to crop text regions for subsequent recognition. These models, however, were limited to handling horizontally oriented text. Extensions of these methods introduced adaptations to manage text with arbitrary orientations and shapes. For example, He et al. Xing et al. (2019) expanded the RoIAlign method to extract features of text in arbitrary orientations, and Sun et al. Liu et al. (2018) introduced a perspective RoI transform layer to align quadrangle proposals into feature maps. More recent innovations, such as the Mask TextSpotter series Liao et al. (2020), have advanced capabilities to handle arbitrarily-shaped text through character-level semantic segmentation, which, while effective, significantly increases the annotation burden due to the necessity for precise character localization.
2.2 Weakly-Supervised Text Spotting
The high cost and labor intensity of annotations required by fully-supervised methods Wang et al. (2025); Feng et al. (2024); Shan et al. (2024); Sun et al. (2024) have spurred interest in weakly-supervised approaches. These methods aim to reduce or entirely eliminate the need for detailed geometric annotations. Notable among the emerging solutions are systems that use minimalistic annotations, such as single-point annotations proposed by SPTS Liu et al. (2023), which represent text positions with single points, reducing annotation complexity. Other exploratory works found in preprints like arXiv suggest innovative directions such as using image captions to train models that can identify text regions without explicit location data. These methods leverage advancements in deep learning, such as contrastive learning and transformer architectures, to intuit text locations from high-level descriptions or minimal guidance. Despite these advancements, there remains a discernible gap between the fully-supervised and weakly-supervised approaches in terms of accuracy and reliability, particularly in complex real-world scenarios. Our work builds on these foundations, aiming to bridge this gap by proposing a transcription-only supervised method that simplifies the annotation process to just text transcriptions, eliminating the need for any geometric data. This approach not only aligns with the trend toward reducing annotation dependencies but also enhances the scalability and accessibility of text spotting technologies.
3 Methodology
The methodology of our EchoSpot introduces several innovative components designed to effectively spot text without reliance on location annotations. The approach leverages the interaction between text queries and image embeddings, refined through sophisticated neural network mechanisms. This section outlines the core components of our methodology.
3.1 Query-based Cross Attention
At the heart of EchoSpot is the Query-based Cross Attention module, which utilizes a transformer architecture similar to those found in natural language processing. The module begins with randomly initialized text queries which are processed through stacked layers of self-attention and cross-attention mechanisms. These mechanisms enable the model to generate embeddings that represent potential text locations based on contextual clues from the image, effectively allowing the model to ”learn” where texts are likely to be located without explicit geometrical annotations.
3.2 Coarse-to-Fine Cross Attention Localization Mechanism
Building on the foundations laid by the initial query-based attention, the Coarse-to-Fine Cross Attention Localization Mechanism refines the text spotting process further. This two-pronged approach first filters through the rough text regions identified by the initial module to discard irrelevant features. It then employs a more refined cross-attention process that sharpens the focus on text regions, enhancing the model’s accuracy in spotting text locations. Text-based Hungarian Matching Loss This component introduces a novel loss function tailored for our transcription-only supervision approach. It uses the Hungarian matching algorithm to align predicted text sequences with their corresponding ground truths during training. This alignment helps in backpropagating accurate error gradients that improve both the recognition and localization aspects of the model, thereby enhancing overall learning efficiency.
3.3 Circular Curriculum Learning Strategy
To address the challenges associated with training from weak annotations, we devise a Circular Curriculum Learning Strategy. This strategy schedules the training process to initially focus on simpler, clearer text instances before gradually introducing more complex images. This pedagogical approach helps the model develop a robust understanding of various text forms and contexts, thereby improving its generalizability and effectiveness on real-world data.
3.4 Audio Annotation
In an effort to simplify the annotation process further and make it accessible, we introduce an audio annotation system. This system allows annotators to speak the text present in images, which is then converted into textual annotations through speech recognition technologies. This method not only speeds up the annotation process but also opens up possibilities for persons with visual impairments to contribute to dataset creation.
3.5 Inference Process
During inference, the model employs the learned embeddings to predict text locations and content. The process begins with extracting features from the input image, followed by the application of the coarse-to-fine mechanism to pinpoint text regions. The predicted regions are then processed for text recognition and classification, providing the final spotting results. Through these methodological innovations, EchoSpot advances the state of text spotting by reducing dependency on costly annotations while maintaining competitive accuracy and efficiency.
4 Experiments
To evaluate the effectiveness of the EchoSpot, we conducted comprehensive experiments across various benchmarks. This section describes the datasets used, the evaluation protocols, implementation details, and a detailed discussion of the results, including an ablation study to validate specific components of our methodology.
4.1 Datasets
Our experiments were performed using several well-known scene text benchmarks:
ICDAR 2013 (IC13) Karatzas et al. (2013): Primarily contains horizontally aligned text.
ICDAR 2015 (IC15)Karatzas et al. (2015): Includes more challenging, multi-oriented text.
Total-Text: Features curved and highly irregular text shapes.
SCUT-CTW1500 Yuliang et al. (2017): Contains curved and annotated text lines.
For training, apart from specific benchmark datasets, we utilized SynthText and COCO-Text to provide a diverse range of text instances, enhancing the model’s ability to generalize across different text types and backgrounds.
4.2 Evaluation Protocol
We adopted two main metrics for evaluation: Polygon Metric: Assesses the precision of the detected text regions against ground truth polygons, using Intersection over Union (IoU) as a measure. Single-point Metric: Evaluates the accuracy of the model in predicting the central point of text instances, beneficial for comparing with methods that do not provide exact bounding boxes.
4.3 Implementation Details
The EchoSpot model was implemented using a modular architecture comprising several key components detailed in the Methodology section. We employed a ResNet-50 backbone augmented with additional convolutional blocks to extract robust features from input images. The model was trained using Adam optimizer with a learning rate strategy adapted to the circular curriculum learning approach, gradually decreasing as the training progressed through cycles of increasing complexity.
4.4 Experimental Results and Ablation Study
Our results demonstrate competitive performance across all benchmarks, with particularly strong results in datasets involving curved and irregular text, underscoring the model’s ability to handle complex text spotting scenarios. The ablation study specifically highlighted the impact of the coarse-to-fine mechanism and the circular curriculum learning strategy. Removing the coarse-to-fine mechanism led to a noticeable drop in performance, validating its role in refining text localization. Similarly, replacing the circular curriculum learning with traditional training approaches resulted in slower convergence and reduced final accuracy, emphasizing its effectiveness in managing the learning process from simple to complex scenes.
4.5 Discussion
The experiments confirm that EchoSpot can achieve high levels of accuracy without relying on detailed geometric annotations, significantly reducing the annotation burden and costs associated with traditional text spotting methods. The model’s robustness across diverse types of text and its ability to learn from minimal supervision suggest promising directions for further research and practical applications in optical character recognition and related fields.
Figure 3: Qualitative results. Images are selected from ICDAR 2013 (first col.), SCUT-CTW1500 (second col.), Total-Text (third col.), and ICDAR 2015 (fourth col.). The first row contains visualizations of single-point, while the second row contains visualizations of masks. As shown in the figure, our method is robust against various text types, including long text, large text, small text, curved text, perspective text, and fuzzy text.
Number of Total-Text End-to-End
Coarse-to-Fine Attention None Full
0 58.5 67.4
1 65.1 74.8
2 66.3 75.2
Table 1: End-to-end recognition results on Total-Text w.r.t number of times of coarse-to-fine cross attention localization mechanism. “None” denotes lexicon-free. “Full” denotes that we use all the words appearing in the test set. All the results are obtained under the single-point metric.
Method IC13 End-to-End
S W G
Fully-supervised methods
FOTS Liu et al. (2018) 88.8 87.1 80.8
TextNet Sun et al. (2018) 89.8 88.9 83.0
Mask TextSpotter Lyu et al. (2018) 93.3 91.3 88.2
Boundary Wang et al. (2020) 88.2 87.7 84.1
Text Perceptron Qiao et al. (2020) 91.4 90.7 85.8
Point-based methods
SPTS (Single-Point) Peng et al. (2021) 87.6 85.6 82.9
Transcription-only methods
EchoSpot (Single-Point) (Ours) 86.4 85.1 82.2
EchoSpot (Polygon) (Ours) 77.7 76.8 73.3
Table 2: End-to-end recognition results on ICDAR 2013. “S”, “W”, and “G” denote recognition with “Strong”, “Weak”, and “Generic” lexicon, respectively. “Single Point” and “Polygon” denote the two metrics mentioned in Sec. LABEL:sec-metric.
Method IC15 End-to-End
S W G
Fully-Supervised methods
FOTS Liu et al. (2018) 81.1 75.9 60.8
Mask TextSpotter Lyu et al. (2018) 83.0 77.7 73.5
CharNet Xing et al. (2019) 83.1 79.2 69.1
TextDragon Feng et al. (2019) 82.5 78.3 65.2
Mask TextSpotter v3 Liao et al. (2020) 83.3 78.1 74.2
MANGO Qiao et al. (2021) 81.8 78.9 67.3
ABCNetV2 Liu et al. (2020b) 82.7 78.5 73.0
PAN++ Wang et al. (2021b) 82.7 78.2 69.2
Point-based methods
SPTS (Single-Point) Peng et al. (2021) 64.6 58.8 54.9
Transcription-only methods
EchoSpot (Single-Point) (Ours) 65.9 59.6 52.4
EchoSpot (Polygon) (Ours) 60.2 54.5 47.1
Table 3: End-to-end recognition results on ICDAR 2015.
Method Total-Text End-to-End
None Full
Fully-Supervised methods
CharNet Xing et al. (2019) 66.2 -
ABCNet Liu et al. (2020a) 64.2 75.7
PGNet Wang et al. (2021a) 63.1 -
Mask TextSpotter Lyu et al. (2018) 65.3 77.4
Qin et al Qin et al. (2019) 67.8 -
Mask TextSpotter v3 Liao et al. (2020) 71.2 78.4
MANGO Qiao et al. (2021) 72.9 83.6
PAN++ Wang et al. (2021b) 68.6 78.6
ABCNet v2 2021ABCNet 70.4 78.1
Point-based methods
SPTS (Single-Point) Peng et al. (2021) 67.9 74.1
Transcription-only methods
EchoSpot (Single-Point) (Ours) 65.1 74.8
EchoSpot (Polygon) (Ours) 61.5 73.0
Table 4: End-to-end recognition results on Total-Text. “None” denotes lexicon-free. “Full” denotes that we use all the words appearing in the test set.
Method SCUT-CTW1500 End-to-End
None Full
Fully-Supervised methods
TextDragon Feng et al. (2019) 39.7 72.4
ABCNet Liu et al. (2020a) 45.2 74.1
MANGO Qiao et al. (2021) 58.9 78.7
ABCNet v2 2021ABCNet 57.5 77.2
Point-based methods
SPTS (Single-Point) Peng et al. (2021) 56.3 67.2
Transcription-only methods
EchoSpot (Single-Point) (Ours) 54.2 65.3
EchoSpot (Polygon) (Ours) 51.4 61.7
Table 5: End-to-end recognition results on SCUT-CTW1500.
Annotation Style Total-Text End-to-End
None Full
Text Transcription 65.1 74.8
Audio (Word) 64.3 73.9
Audio (Character) 64.9 74.5
Table 6: End-to-end recognition results w.r.t annotation styles on Total-Text. “Word” and “Character” denote word-by-word and character-by-character pronunciation, respectively.
5 Conclusion
In this paper, we introduced the EchoSpot, a novel framework designed to address the challenges of scene text spotting with minimal reliance on costly and labor-intensive geometric annotations. By leveraging a query-based interaction between text queries and image embeddings, EchoSpot efficiently learns to spot text locations and recognize text content using only transcription annotations. Our approach significantly simplifies the annotation process by allowing for text annotations through transcription and even audio inputs, making it not only cost-effective but also accessible to a broader range of contributors, including those with disabilities. The introduction of a circular curriculum learning strategy ensures that the model can be effectively trained from scratch on weakly annotated data, gradually learning to handle increasingly complex scenes through a structured learning approach. The experimental results across several challenging datasets demonstrate that EchoSpot achieves competitive performance compared to state-of-the-art methods that rely on full supervision. Particularly, EchoSpot performs well on benchmarks involving curved and arbitrary-shaped text, underscoring its robustness and adaptability. Future work will explore further enhancements to the coarse-to-fine localization mechanism to improve the precision of text localization. Additionally, extending the framework to support more languages and script types will increase its applicability on a global scale. We also see potential in exploring deeper integrations of audio-visual data to enrich the training process and improve the model’s performance in multimodal contexts. Overall, the innovations presented in EchoSpot pave the way for more efficient and scalable text-spotting solutions, promising significant advancements in how optical character recognition technology is applied across various real-world applications.

### v3

@@
 (2024b); feng2023docpedia; Feng et al.
-(2023) has evolved from focusing primarily on horizontal or multi-oriented text to addressing more complex arbitrarily-shaped text.
+(2023) has evolved from focusing primarily on horizontal or multi-oriented text to addressing more complex, arbitrarily-shaped text.
 This evolution has prompted shifts from traditional annotation formats like horizontal rectangles and quadrilaterals to more sophisticated polygonal annotations, as depicted in Figure 1.
-Traditionally, scene text spotting involved separate processes for text detection and recognition, where both tasks required annotations for training—text recognition used transcription annotations, and text detection depended on geometric annotations like polygons or bounding boxes.
+Traditionally, scene text spotting involved separate processes for text detection and recognition, where both tasks required annotations for training.
+Text recognition used transcription annotations, and text detection depended on geometric annotations like polygons or bounding boxes.
 The reliance on detailed geometric annotations for training not only escalates the cost but also increases the manpower required, making the process less efficient and scalable.
@@
 Given these challenges, there is a growing interest in exploring methods that reduce reliance on expensive annotations.
-Preliminary efforts, such as the single-point annotation strategy proposed by SPTS Liu et al.
+Preliminary efforts Wang et al.
+(2024, 2025b); Sun et al.
+(2025), such as the single-point annotation strategy proposed by SPTS Liu et al.
 (2023), have marked a significant step forward but still leave room for improvement.
@@
 The high cost and labor intensity of annotations required by fully-supervised methods Wang et al.
-(2025); Feng et al.
+(2025a); Feng et al.
 (2024); Shan et al.